OpenStudy (anonymous):

here is the problem: f(x)=x^2-4x-3. I have the x cordinate, which is 2. Can someone please explain to me how y=-7?

OpenStudy (anonymous):

f(x) = x^2 - 4x - 3 Since you know that x = 2, you just need to plug 2 into the equation wherever you see x! With that said: f(2) = 2^2 - (2*4) - 3 = 4 - 8 - 3 = -4 -3 = -7

OpenStudy (anonymous):

Thanks